OpenStudy (anonymous):

If H(x)=sqrt(x+12) Find H'(x)

myininaya (myininaya):

\[H(x)=f(g(x))=> H'(x)=f'(g(x))g'(x)\] where g(x)=x+12 and f(x)=x^(1/2)

OpenStudy (anonymous):

so what would the final answer be? sorry, i havent gotten this stuff down yet. need as much help as possible.

myininaya (myininaya):

use what i gave you to find H' g'(x)=1 f'(x)=1/2 * x^(-1/2) no plug g(x) into f'(x) yo fin f'(g(x) and then don't forget to multiply that by g'(x)=1

OpenStudy (anonymous):

why is g'(x)=1 shouldn't it be x?

OpenStudy (anonymous):

?

myininaya (myininaya):

because g=x+12 g'=1 since the slope of x+12 is 1

OpenStudy (anonymous):

and after i do what you said i get 1/2(x+12^(-1/2)) but thats not the final answer according to my hw website

myininaya (myininaya):

no it is not you have to simplify you need to close a parantheis too

myininaya (myininaya):

you put the paranthesis in the wrong spot

myininaya (myininaya):

after you fix that write your answer so that any exponents are positve

myininaya (myininaya):

\[\frac{1}{2}(x+12)^\frac{-1}{2}=\frac{1}{2} \cdot \frac{1}{(x+12)^\frac{1}{2}}\] \[\frac{1}{2\sqrt{x+12}}\]

OpenStudy (anonymous):

i seem to be putting it in wrong. how would i put it in the old school way?

myininaya (myininaya):

1/[2(x+12)^(1/2)]

OpenStudy (anonymous):

ok thanks. sorry for all the trouble.